Calculate the value of the translational partition function of O2 at 300 K for a volume of 1 m3. Submit your answer multiplied by 10−32 (i.e. if you calculate 4.2e33, input 42).
1
Expert's answer
2020-06-03T13:34:05-0400

qT = V / Λ3 = V / (h / (2πmkBT)1/2)3 = 1 / (6.626 × 10-34 / (2 × 3.142 × (0.032 / (6.022 × 1023) × 1.38 × 10-23 × 300))1/2)3 = 2.78 × 1033

Answer: 27.8.